# Eduard Alarcón1

### v2

@@ I Introduction
Nevertheless, the recent advancement of fabricating cryo-CMOS electronics has paved the way of integrating the control electronics at 4 K, with ultra-low power dissipation [7, 8]. With these advancements, RF communication inside the quantum structure makes a compelling case for multi-core quantum systems, with the necessity of global connectivity and reduced latency despite of the distance, while proving reconfigurable links with broadcasting capabilities. In addition, with the integrated RF on-chip antenna and transmitter and receiver circuits at 4 K, the thermal noise added in cryogenic temperature will be low, while increasing the Signal to Noise Ratio (SNR) of the wireless link.
@@ II-C Environment Description
Fig. 2 illustrates the structure of the simulation model with integrated on-chip antennas. As described in Section II-B, the designed differential dipole antenna for cryogenic temperatures was integrated into the cryostat structure by placing it on the printed circuit board (PCB), which is positioned above the cooling plate that is connected to the central cooling tube. The dimensions of the cryostat are based on a standard design used for quantum computing experiments in the ranges of, diameter d=30 cm and height, h=70 cm [16]. The distance between the bottom-middle and middle-top cooling plates are 10 cm and 15 cm, respectively. As the material and structural properties were effecting the antenna resonance, the antennas are placed on top of the cooling plates with a 6 cm gap for effective performance.
@@ III-B Cryo-CMOS antenna within the Cryostat Scenario
To ensure proper impedance matching and efficient power transfer, the free-space dipole optimal length ($L_{opt}$) of 2.8 mm was replaced with 3.06 mm. In addition, the length of the SiO2 and Si layers ($L_{layers}$), as well as the thickness of the substrate, were explored. An optimal resonance was obtained with $L_{layers}$ of 3.72 mm and Si thickness of 0.3 mm. As depicted in Fig. 3(a), the on-chip dipole resonates at around 28 GHz with an approximate value of -21 dB. The efficiency and realized gain of the antenna were also assessed inside the cryostat. Fig. 3(b) shows an efficiency of 77%, which reflects the effects of the cryostat environment on the antenna resonance. In addition, a realized gain of 8.25 dBi was obtained at 28 GHz.
The EM field distribution of the on-chip antenna was also simulated. Fig. 4(a) depicts the signal propagating from the GSSG pads to the dipole arms, the coupling and signal propagation around the dipoles, where the differential feeding introduces a 180° phase shift between them. Additionally, EM field distribution inside the cryostat model is observed in Fig. 4(b). It is shown that the electric fields are oscillating around the enclosed chamber with multi-path signal propagation. However as the QPUs are place one temperature level below, the effect of interference is reduced.
@@ IV Conclusion
In this paper, we propose and evaluate an on-chip differential dipole antenna designed for operation at 28 GHz within a quantum computer structure. By leveraging cryo-CMOS technology and material properties adapted for cryogenic environments, the dipole exhibits enhanced performance compared to room temperature, achieving reflection coefficients of -28.85 dB in free space with and -21 dB inside the cryostat corresponding to total efficiencies of 92% and 77% respectively. Furthermore, the simulations demonstrate the impact of substrate thickness, layer dimensions on the antenna performance while initially assessing the EM distribution inside a realistic cryostat computer.

### v3

@@ I Introduction
Nevertheless, the recent advancement of fabricating cryo-CMOS electronics has paved the way of integrating the control electronics at 4 K, with ultra-low power dissipation [7, 8]. With these advancements, RF communication inside the quantum structure makes a compelling case for multi-core quantum systems, with the necessity of global connectivity and reduced latency despite of the distance, while proving reconfigurable links with broadcasting capabilities. In addition, with the integrated RF on-chip antenna and transceiver circuits at 4 K, the thermal noise added in cryogenic temperature will be low, while increasing the Signal to Noise Ratio (SNR) of the wireless link.
@@ II-C Environment Description
Fig. 2 illustrates the structure of the simulation model with integrated on-chip antennas. As described in Section II-B, the designed differential dipole antenna for cryogenic temperatures was integrated into the cryostat structure by placing it on the printed circuit board (PCB), which is positioned above the cooling plate that is connected to the central cooling tube. The dimensions of the cryostat are based on a standard design used for quantum computing experiments in the ranges of, diameter d=30 cm and height, h=70 cm [16]. The distance between the bottom-middle and middle-top cooling plates are 10 cm and 15 cm, respectively.
@@ III-B Cryo-CMOS antenna within the Cryostat Scenario
To ensure proper impedance matching and efficient power transfer, the free-space dipole optimal length ($L_{opt}$) of 2.8 mm was replaced with 3.06 mm. In addition, the length of the SiO2 and Si layers ($L_{layers}$), as well as the thickness of the substrate, were explored. An optimal resonance was obtained with $L_{layers}$ of 3.72 mm and Si thickness of 0.3 mm. As depicted in Fig. 3(a), the on-chip dipole resonates at around 28 GHz with an approximate value of -21 dB. The efficiency and realized gain of the antenna were also assessed inside the cryostat. Fig. 3(b) shows an efficiency of 77%, which reflects the effects of the cryostat environment on the antenna resonance. In addition, a realized gain of 8.25 dBi was obtained at 28 GHz. However, thermal asymmetries near the differential feed lines can trigger a common-mode conversion, which have to be carefully addressed in cryogenic conditions.
The EM field distribution of the on-chip antenna was also simulated. Fig. 4(a) depicts the signal propagating from the GSSG pads to the dipole arms, the coupling and signal propagation around the dipoles, where the differential feeding introduces a 180° phase shift between them. Additionally, EM field distribution inside the cryostat model is observed in Fig. 4(b). It is shown that the electric fields are oscillating around the enclosed chamber with multi-path signal propagation. However, as the QPUs are placed one temperature level below, the impact of radiated power is reduced.
@@ IV Conclusion
In this paper, we propose and evaluate an on-chip differential dipole antenna designed for operation at 28 GHz within a quantum computer structure. By leveraging cryo-CMOS technology and material properties adapted for cryogenic environments, the dipole exhibits enhanced performance compared to room temperature, achieving reflection coefficients of -28.85 dB in free space with and -21 dB inside the cryostat corresponding to total efficiencies of 92% and 77% respectively. The simulations further demonstrate the impact of substrate thickness, layer dimensions on the antenna performance while initially assessing the EM distribution inside a realistic cryostat computer.